# Unveiling the Blind Spots: A Critical Examination of Fairness in Autonomous Driving Systems

### v2

@@ 3.4. Evaluation Metric
We employ Statistical Parity Difference (SPD) and Equal Opportunity Difference (EOD), which are the most widely-adopted fairness metrics in the literature (Hort et al., 2023). SPD quantifies the difference in probabilities of favorable outcomes between unprivileged and privileged groups; EOD indicates the difference in the true-positive rate between unprivileged and privileged groups. In the context of pedestrian detection, both SPD and EOD measure the disparity in proportions of successfully detected pedestrians between privileged and unprivileged groups. They also both express the difference in miss rates between privileged and unprivileged groups.
Miss Rate (MR) is the most commonly-studied performance metric in pedestrian detection (Brandao, 2019; Kogure et al., 2022), which quantifies the proportion of undetected pedestrians. Formally, it is calculated as follows:
@@ 3.5. Statistical Analysis
To assess the extent to which any observed unfairness is statistically significant (i.e., whether there is a significant difference in the miss rate between privileged and unprivileged groups) we use the two-proportion z-test (Casella and Berger, 2007). This statistical test is widely adopted to analyze differences between two proportions (Niu et al., 2016; Veizaga et al., 2021). A result is deemed significant only if the obtained $p$-value falls below 0.05.
@@ 4.1. RQ1: Overall Fairness
RQ1 investigates the overall fairness of eight state-of-the-art pedestrian detectors with respect to gender, age, and skin tone. First, for each detector, we compute the miss rate (MR) for different demographic groups and calculate EOD based on the MR results over all the datasets that we study. We also use the two-proportion z-test to determine the significance of any observed unfairness (as described in Section 3.5). Table 6 presents the results, with significant unfairness results highlighted in shading. Next, we analyze the results for gender, age, and skin tone, respectively.
Table 6. RQ1: Overall fairness with respect to gender, age and skin tone. Statistically significant unfairness (i.e., MR difference) is highlighted in shading.
@@
Then, we explore the difference across the four datasets used in our study. We find that in the CityPersons and EuroCity-Day datasets, which exclusively contain day-time data, there are no significant differences in the miss rate between females and males. However, the results on the EuroCity-Night dataset present a contrasting observation, where all detectors demonstrate a higher miss rate for females, thus illustrating bias in detecting females. In the BDD100k dataset that includes both day-time and night-time images, the performance difference is less pronounced. These observations motivate us to hypothesize that brightness conditions may influence the fairness of pedestrian detectors, which is further investigated in RQ2.
@@
Furthermore, we illustrate the miss rate difference across the four datasets that we use in Figure 1, with significant age unfairness indicated by labeled miss rate values. From the figure, we observe that the miss rate of children is consistently higher than that of adults for all the datasets and all detectors. In particular, out of the total 32 results (combinations of four datasets and eight pedestrian detection models), 30 exhibit a statistically significant difference in the miss rate with $p$-value ¡ 0.05. This indicates that the bias favoring adults is not specific to a particular dataset or model, highlighting a strong unfairness between adults and children.
Figure 1. RQ1: Miss rates for children and adults in each dataset. Statistically significant age unfairness (i.e., miss rate difference) is indicated by labeled miss rate values.
The age bias can be attributed to the inherent challenge of detecting small objects in object and person detection research, owing to the limited information provided by small bounding box sizes (Zhang et al., 2016; Dollar et al., 2009; Kogure et al., 2022; Cao et al., 2020). Given that children generally have smaller bodies compared to adults, their bounding boxes in the images also tend to be smaller. To verify this, we analyze the distribution of bounding box sizes for pedestrians detected and undetected by all pedestrian detectors, as well as the distribution of ground-truth bounding box sizes for both adults and children. The results, presented in Figure 2, reveal a correspondence between the two distributions. Children as well as undetected pedestrians tend to have smaller bounding boxes.
Figure 2. Bounding boxes distribution of adults and children (left) and bounding boxes distribution of undetected and detected pedestrians (right).
Skin tone. As shown in Table 6, current pedestrian detectors indicate an impartial performance between light-skin and dark-skin individuals, with a minor average miss rate discrepancy of 0.44% between light-skin and dark-skin groups. Within the scope of eight detectors assessed, three were observed to have marginally increased miss rates for the light-skin group relative to the dark-skin group (with three of them being statistically significant).
Ans. to RQ1: On average, the pedestrian detectors currently in use do not show significant performance differences when considering gender and skin tone. However, a pronounced disparity is observed across different age groups. Specifically, the data points to a substantial miss rate of 20.14% when detecting between children and adults. This significant margin underscores a heightened risk to child pedestrians within the ambit of present autonomous driving systems.
@@ 4.2. RQ2: Fairness in Different Scenarios
RQ2 investigates the fairness of state-of-the-art pedestrian detectors under different real-world autonomous driving scenarios.
@@
As introduced in Section 3.3.3, we consider brightness conditions from two perspectives: day time and night time.
@@
We first evaluate the average miss rates of the eight pedestrian detectors under day time and night time. The results, presented in Table 7, indicate a noticeable increase in average miss rates during night time compared to day time. For example, the average miss rates for males and females at night are 33.20% and 36.57%, respectively, compared to 17.24% and 17.32% during the day. A similar pattern is observed for age and skin tone. This indicates that the transition from day time to night time influences the performance of pedestrian detectors, with statistically significant higher miss rates observed at night.
Table 7. RQ2.1: MR and EOD of each pedestrian detector under day time and night time. The statistically significant unfairness results are highlighted in shading.
Detectors Gender Age Skin Tone (LS: Light Skin, DS: Dark Skin)
Day-time Night-time Day-time Night-time Day-time Night-time
MR Male MR Female EOD MR Male MR Female EOD MR Adult MR Child EOD MR Adult MR Child EOD MR LS MR DS EOD MR LS MR DS EOD
YOLOX 9.73% 9.38% 0.35% 9.99% 14.53% -4.54% 10.80% 40.77% -29.97% 17.81% 54.93% -37.12% 4.89% 4.03% 0.86% 7.06% 2.52% 4.53%
RetinaNet 10.42% 11.79% -1.37% 11.99% 15.48% -3.49% 12.73% 41.92% -29.19% 18.97% 61.97% -43.01% 7.44% 4.33% 3.11% 13.38% 5.04% 8.34%
Faster RCNN 3.66% 3.59% 0.07% 4.41% 6.06% -1.66% 4.50% 25.00% -20.50% 7.34% 33.80% -26.46% 5.36% 3.13% 2.23% 9.00% 4.20% 4.80%
Cascade RCNN 3.58% 3.68% -0.10% 5.11% 5.87% -0.76% 4.52% 25.58% -21.05% 6.81% 33.80% -26.99% 4.84% 2.99% 1.86% 7.30% 3.36% 3.94%
ALFNet 24.52% 24.39% 0.13% 57.29% 63.42% -6.14% 26.52% 49.42% -22.90% 65.10% 83.10% -18.00% 37.05% 38.66% -1.61% 73.48% 73.95% -0.47%
CSP 26.76% 26.38% 0.38% 61.57% 67.09% -5.51% 28.64% 46.92% -18.28% 63.42% 76.06% -12.64% 55.90% 59.85% -3.95% 92.21% 92.44% -0.22%
MGAN 25.48% 26.29% -0.81% 47.59% 47.76% -0.17% 27.87% 43.46% -15.59% 49.72% 69.01% -19.30% 47.12% 49.70% -2.58% 85.64% 84.03% 1.61%
PRNet 33.73% 33.07% 0.66% 67.63% 72.33% -4.70% 34.80% 56.92% -22.13% 72.60% 92.96% -20.36% 54.39% 54.63% -0.24% 89.29% 86.55% 2.74%
Average 17.24% 17.32% -0.09% 33.20% 36.57% -3.37% 18.80% 41.25% -22.45% 37.72% 63.20% -25.48% 27.31% 27.16% 0.15% 47.17% 44.01% 3.16%
@@
Then, we investigate whether the performance is equally decreased for different demographic groups. Specifically, we explore the fairness change from day time to night time. Table 7 shows the results, with statistically significant unfairness (i.e., miss rate difference) emphasized in shading.
@@
Regarding gender, we observe a shift in the miss rate difference between males and females from day time to night time. During the day, there is only a slight -0.09% difference for males and females. However, during night time, the difference increases to -3.37% with statistically significant, indicating a notable change in fairness. In the night-time condition, all pedestrian detectors exhibit higher miss rates for females compared to males, with six of them showing statistically significant differences.
@@
For age, the miss rate difference for children and adults increases from day time to night time, with the average difference increasing from -22.45% during the day to -25.48% at night with statistically significant. This suggests a higher probability of children being undetected during night time conditions.
@@
For skin tone, the miss rate difference between dark-skin and light-skin groups increases from day time to night time, with the average difference increasing from 0.15% at day time to 3.16% at night time. This implies a small increased detection difficulty for the light-skin group at night.
Ans. to RQ2.1: Lower brightness conditions amplify the bias towards children and female in pedestrian detection.
@@
Figure 3. RQ2.2: Average MR of eight pedestrian detectors and EOD trends under different contrast levels.
Following the road map of RQ2.1, we first compare the average miss rates of all pedestrian detectors at different contrast levels, as shown in Figure 3. For skin tone and age, we observe that the average miss rate increases at a lower contrast level (level 1) compared to the other two levels. However, for gender, the average miss rate of all pedestrian detectors does not show an obvious variation across different contrast levels.
Then, we investigate the fairness changes, as indicated by the EOD trends in Figure 3.
For skin tone, as the contrast level decreases, we observe varied responses from pedestrian detectors. Specifically, models like CSP and PRNet exhibit increased difficulty in detecting dark-skin individuals. Conversely, detectors such as RetinaNet and Faster RCNN demonstrate more challenges in identifying the light-skin group. These results demonstrate that contrast conditions have a potential impact on the detection of different skin tone groups
For gender, a reverse trend in EOD emerges. With decreasing contrast levels, pedestrian detectors progressively show a higher miss rate for females than for males. This observation indicates that low contrast conditions negatively impact the detection of females, thereby contributing another factor to the bias and unfairness in pedestrian detection.
For age, the EOD between children and adults maintains a large difference across all contrast levels. Moreover, general object detectors, including YOLOX, RetinaNet, Faster RCNN, and Cascade RCNN, exhibit increased miss rates difference at lower contrast levels. This observation implies that contrast levels in general object detectors could potentially exacerbate the bias, amplifying the disparity in detection performance between adults and children.
Ans. to RQ2.2: Low-contrast conditions have a negative influence on the fairness of some existing pedestrian detectors regarding gender, skin tone and age, especially posing a greater challenge in detecting children and females.
@@
Table 8 presents the miss rates of the eight pedestrian detectors under non-rainy and rainy weather conditions, with statistically significant unfairness results shaded.

### v3

@@ 3.4. Evaluation Metric
There have been well-established quantitative measures for group fairness in the literature. The most widely-adopted fairness measures include SPD (Statistical Parity Difference), EOD (Equal Opportunity Difference), and AOD (Average Odds Difference) (Chen et al., 2022, 2024; Zhang and Harman, 2021). SPD quantifies the difference in probabilities of favorable outcomes between unprivileged and privileged groups; EOD indicates the difference in the true-positive rate between unprivileged and privileged groups; AOD calculates the average difference between the false-positive rate and true-positive rate for unprivileged and privileged groups.
In the context of pedestrian detection, both SPD and EOD measure the disparity in proportions of successfully detected pedestrians between privileged and unprivileged groups. They also both express the difference in miss rates between privileged and unprivileged groups. Miss Rate (MR) is the most commonly-studied performance metric in pedestrian detection (Brandao, 2019; Kogure et al., 2022), which quantifies the proportion of undetected pedestrians. Formally, it is calculated as follows:
@@
The calculation of AOD requires precise false-positive information, referring to instances where members of the negative class (non-pedestrians) are incorrectly classified as the positive class (pedestrians). As described in Section 3.3.2, we adhere to the standard practice in the literature, where we focus on a “reasonable subset” of pedestrians within the images. This approach presents challenges in calculating precise false-positives for each group because the negative class (non-pedestrian) may also include instances that belong to the positive class (pedestrian). Therefore, we do not consider AOD in this study.
In summary, we use both SPD and EOD as fairness measures for evaluating pedestrian detectors. Since these two measures yield identical values for pedestrian detection, for the remainder of the paper, we present only EOD.
@@ 3.5. Statistical Analysis
To assess the extent to which any observed unfairness is statistically significant (i.e., whether there is a significant difference in the miss rate between privileged and unprivileged groups), we use the two-proportion z-test (Casella and Berger, 2007). This statistical test is widely used to analyze differences between proportions (Niu et al., 2016; Veizaga et al., 2021). A result is deemed significant only if if the obtained $p$-value falls below a predetermined threshold (in our case, 0.05, a widely-accepted threshold in the fairness literature (Chen et al., 2022, 2023)). For instance, in evaluating whether there exists a difference between the miss rates for male ($MR_{male}$) and female ($MR_{female}$) individuals detected by a pedestrian detector, the null hypothesis assumes that $MR_{male}$ is equal to $MR_{female}$. If the resulting $p$-value is lower than 0.05, we reject the null hypothesis, indicating a significant difference between $MR_{male}$ and $MR_{female}$.
3.6. Experimental details
The experiments are implemented based on open-source frameworks of each pedestrian detector. For general object detectors, we select the pre-trained models with the highest accuracy from the the MMdetection model zoo  (MMDetection Developers, nd). For pedestrian-specific detectors, we employ pre-trained models available from their respective public repositories  (hasanirtiza, 2021; sxlpris, 2020; liuwei16, 2018).
To ensure the reliability of our results, all experiments are repeated ten times. The final results were derived by calculating the average across these ten iterations.
All experiments are performed on a platform equipped with 64GB RAM, 2.5GHz Intel Xeon (R) v3 Dual CPUs, and one NVIDIA GeForce RTX 2080 Ti GPU. YOLOX, RetinaNet, Faster RCNN, and Cascade RCNN are implemented using PyTorch 1.8.1 and Python 3.7 on Ubuntu 18.04 LTS, following the MMdetection configuration (Chen et al., 2019). CSP and MGAN use PyTorch 1.10.0 and Python 3.8 on Ubuntu 18.04 LTS, adhering to the Pedestron configuration (Hasan et al., 2022). ALFNet (Liu et al., 2018) and PRNet (Song et al., 2020) are implemented using Keras 2.0.6, Tensorflow 1.4.0, and Python 2.7 on Ubuntu 16.04 LTS.
@@ 4.1. RQ1: Overall Fairness
RQ1 investigates the overall fairness of eight state-of-the-art pedestrian detectors regarding gender, age, and skin tone. First, for each detector, we compute the miss rate (MR) for different demographic groups and calculate EOD based on the MR results over all the datasets that we study. We also use the two-proportion z-test to determine the significance of any observed unfairness, as described in Section 3.5. Table 6 presents the results, with significant unfairness results highlighted in shading. In the following, we analyze the results for gender, age, and skin tone, respectively.
Table 6. (RQ1) Overall fairness in pedestrian detection across gender, age, and skin tone. Statistically significant biases, indicated by EOD (i.e., miss rate difference), are shaded. On average, detectors display comparable miss rates for female and male pedestrians, as well as for dark-skin and light-skin individuals. However, concerning age, significant bias is observed, as pedestrian detectors exhibit a 20.14% higher miss rate for children compared to adults.
@@
Furthermore, we analyze the miss rate difference achieved by each pedestrian detector across the four datasets used in our study. Figure 2 illustrates the results. Significant gender biases (i.e., significant miss rate differences) are indicated by labeled miss rate values. We observe that in the CityPersons and EuroCity-Day datasets, containing only day-time data, only one detector in the EuroCity-Day datasets exhibits significant differences in miss rates between females and males. For the remaining results, there are no notable differences in miss rates between genders. However, results from the EuroCity-Night dataset present a contrasting observation, where seven of eight detectors exhibit a significantly higher miss rate for females, indicating bias in detecting females. In the BDD100k dataset that includes both day-time and night-time images, the miss rate difference is less pronounced. These observations motivate us to hypothesize that brightness conditions may influence the fairness of pedestrian detectors, which is further investigated in RQ2.
Figure 2. (RQ1) Miss rates of pedestrian detectors for females and males across datasets. Statistically significant gender biases are indicated by labeled miss rate values. In CityPersons and EuroCity-Day datasets with only day-time data, no significant gender biases are observed. However, in the EuroCity-Night dataset, seven out of eight detectors show significantly higher miss rates for females, revealing bias in female detection.
@@
Furthermore, we illustrate the miss rate difference across four datasets in Figure 3, with significant age biases indicated by labeled miss rate values. We observe that the miss rate of children is consistently higher than that of adults across all the datasets and all detectors. In particular, out of the total 32 results (combinations of four datasets and eight pedestrian detection models), 30 exhibit a statistically significant miss rate difference (i.e., $p$-value ¡ 0.05). This indicates that the bias favoring adults is not specific to a particular dataset or detector, highlighting a strong unfairness between adults and children.
Figure 3. (RQ1) Miss rates of pedestrian detectors for children and adults across datasets. Statistically significant age biases are labeled with miss rate values. In 30 out of 32 scenarios (comprising four datasets and eight detectors), children have significantly higher miss rates than adults.
The age bias may be attributed to the inherent challenge of detecting small objects, owing to the limited information provided by small bounding boxes (Zhang et al., 2016; Dollar et al., 2009; Kogure et al., 2022; Cao et al., 2020). Given that children generally have smaller bodies compared to adults, their bounding boxes in the images also tend to be smaller. To demonstrate this, we analyze the distribution of bounding box sizes for pedestrians detected and undetected by all pedestrian detectors, as well as the distribution of ground-truth bounding box sizes for both adults and children. The results, presented in Figure 4, reveal a correspondence between the two distributions. Children as well as undetected pedestrians tend to have smaller bounding boxes.
Figure 4. (RQ1) Bounding box size distributions of adults and children (left) and bounding box size distributions of undetected and detected pedestrians (right). We observe that both children and undetected pedestrians tend to have smaller bounding boxes.
Skin tone. As displayed in Table 6, current pedestrian detectors exhibit minimal bias between light-skin and dark-skin individuals, with an average miss rate difference of just 0.44%. Specifically, among the four pedestrian-specific detectors, none show significant skin tone bias, whereas among the four general object detectors, three display significant skin tone bias, with miss rate differences ranging from 2.17% to 3.9%. These disparities are notably smaller than those observed in age bias.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false]
Ans. to RQ1: On average, the eight state-of-the-art pedestrian detectors under study exhibit no significant performance difference in across gender and skin tone, but notable bias across age. In particular, pedestrian detectors exhibit a 20.14% higher miss rate for children compared to adults.
@@ 4.2. RQ2: Fairness in Different Scenarios
RQ2 evaluates the fairness of state-of-the-art pedestrian detectors under different real-world autonomous driving scenarios.
@@
As introduced in Section 3.3.3, we consider two brightness conditions: day time and night time.
@@
We first evaluate the overall miss rates of the eight pedestrian detectors under day time and night time. The results, presented in Table 7, show a noticeable increase in average miss rates during night time compared to day time for each demographic group. For example, the average miss rates for males and females at night are 33.20% and 36.57%, respectively, compared to 17.24% and 17.32% during the day. A similar pattern is observed for age and skin tone. This indicates that the transition from day time to night time influences the performance of pedestrian detectors, with statistically significant higher miss rates observed at night.
@@
Then, we investigate whether the performance is equally decreased for different demographic groups. Specifically, we explore the fairness change from day time to night time. Table 7 shows the results, with statistically significant biases (i.e., miss rate differences) emphasized in shading.
@@
Regarding gender, we observe a shift in the miss rate difference between males and females from day time to night time. During the day, there is only a slight -0.09% difference for males and females. However, during night time, the difference increases to -3.37% with statistically significant, indicating a notable change in fairness. In the nighttime condition, all pedestrian detectors exhibit higher miss rates for females compared to males, with six of them showing statistically significant differences. This potentially worsens concerns regarding female safety during nighttime out, a prevalent societal issue (Wang et al., 2023).
@@
For age, all detectors exhibit significant biases during both day time and night time. Moreover, the miss rate difference for children and adults increases from day time to night time, with the average difference increasing from -22.45% during the day to -25.48% at night with statistically significant. This suggests a higher probability of children being undetected during night time.
@@
For skin tone, the miss rate difference between dark-skin and light-skin groups increases from day time to night time, with the average difference increasing from 0.15% at day time to 3.16% at night time. Nevertheless, the overall skin tone bias is not statistically significant during both day time and night time.
Table 7. (RQ2.1) Miss rates and EOD of each pedestrian detector under day time and night time. Statistically significant unfairness results are shaded. We find that reduced brightness conditions not only decrease the performance of pedestrian detectors but also exacerbate their biases. Notably, while pedestrian detectors generally do not exhibit significant gender bias during day time, biases against females become pronounced with six out of eight detectors showing significant biases during night time.
Gender
Detectors Day-time Night-time
MR Male MR Female EOD (Gender) MR Male MR Female EOD (Gender)
YOLOX 9.73% 9.38% 0.35% 9.99% 14.53% -4.54%
RetinaNet 10.42% 11.79% -1.37% 11.99% 15.48% -3.49%
Faster RCNN 3.66% 3.59% 0.07% 4.41% 6.06% -1.66%
Cascade RCNN 3.58% 3.68% -0.10% 5.11% 5.87% -0.76%
ALFNet 24.52% 24.39% 0.13% 57.29% 63.42% -6.14%
CSP 26.76% 26.38% 0.38% 61.57% 67.09% -5.51%
MGAN 25.48% 26.29% -0.81% 47.59% 47.76% -0.17%
PRNet 33.73% 33.07% 0.66% 67.63% 72.33% -4.70%
Average 17.24% 17.32% -0.09% 33.20% 36.57% -3.37%
Age
Detectors Day-time Night-time
MR Adult MR Child EOD (Age) MR Adult MR Child EOD (Age)
YOLOX 10.80% 40.77% -29.97% 17.81% 54.93% -37.12%
RetinaNet 12.73% 41.92% -29.19% 18.97% 61.97% -43.01%
Faster RCNN 4.50% 25.00% -20.50% 7.34% 33.80% -26.46%
Cascade RCNN 4.52% 25.58% -21.05% 6.81% 33.80% -26.99%
ALFNet 26.52% 49.42% -22.90% 65.10% 83.10% -18.00%
CSP 28.64% 46.92% -18.28% 63.42% 76.06% -12.64%
MGAN 27.87% 43.46% -15.59% 49.72% 69.01% -19.30%
PRNet 34.80% 56.92% -22.13% 72.60% 92.96% -20.36%
Average 18.80% 41.25% -22.45% 37.72% 63.20% -25.48%
Skin Tone (LS: Light Skin, DS: Dark Skin)
Detectors Day-time Night-time
MR LS MR DS EOD (Skin) MR LS MR DS EOD (Skin)
YOLOX 4.89% 4.03% 0.86% 7.06% 2.52% 4.53%
RetinaNet 7.44% 4.33% 3.11% 13.38% 5.04% 8.34%
Faster RCNN 5.36% 3.13% 2.23% 9.00% 4.20% 4.80%
Cascade RCNN 4.84% 2.99% 1.86% 7.30% 3.36% 3.94%
ALFNet 37.05% 38.66% -1.61% 73.48% 73.95% -0.47%
CSP 55.90% 59.85% -3.95% 92.21% 92.44% -0.22%
MGAN 47.12% 49.70% -2.58% 85.64% 84.03% 1.61%
PRNet 54.39% 54.63% -0.24% 89.29% 86.55% 2.74%
Average 27.31% 27.16% 0.15% 47.17% 44.01% 3.16%
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false]
Ans. to RQ2.1: Lower brightness conditions not only diminish the performance of pedestrian detectors but also exacerbate their bias. Particularly, during day time, pedestrian detectors generally do not exhibit significant gender bias, whereas six out of eight detectors demonstrate significant biases against females during night time.
@@
Table 8. (RQ2.2) Miss rates and EOD of each pedestrian detector under different contrast levels. On average, detectors exhibit the most biased results (i.e., the largest absolute value of EOD) under level 2, while demonstrating the fairest outcomes under level 3.
Gender
Detectors level 3 level 2 level 1
Male MR Female MR EOD Male MR Female MR EOD Male MR Female MR EOD
YOLOX 6.30% 5.21% 1.09% 9.42% 10.42% -1.00% 13.99% 14.71% -0.72%
RetinaNet 7.89% 9.46% -1.56% 10.47% 12.76% -2.29% 13.99% 14.24% -0.25%
Faster RCNN 3.83% 2.34% 1.49% 3.66% 4.16% -0.50% 4.34% 5.32% -0.98%
Cascade RCNN 3.43% 2.44% 0.98% 3.65% 4.26% -0.62% 5.12% 5.01% 0.11%
ALFNet 33.33% 31.67% 1.66% 30.28% 32.66% -2.38% 31.09% 34.74% -3.65%
CSP 40.11% 37.41% 2.70% 32.61% 34.80% -2.19% 31.54% 35.45% -3.90%
MGAN 38.52% 38.89% -0.38% 29.25% 31.23% -1.98% 24.61% 24.10% 0.51%
PRNet 48.41% 46.97% 1.43% 39.45% 41.52% -2.06% 36.92% 38.11% -1.19%
Average 22.73% 21.80% 0.93% 19.85% 21.48% -1.63% 20.20% 21.46% -1.26%
Age
Detectors level 3 level 2 level 1
Adult MR Child MR EOD Adult MR Child MR EOD Adult MR Child MR EOD
YOLOX 6.91% 32.89% -25.98% 12.05% 40.92% -28.87% 18.23% 53.23% -35.00%
RetinaNet 10.28% 36.84% -26.56% 13.85% 41.94% -28.10% 18.70% 56.45% -37.75%
Faster RCNN 4.00% 23.68% -19.69% 5.08% 26.09% -21.00% 6.56% 27.42% -20.86%
Cascade RCNN 3.33% 19.74% -16.41% 5.05% 25.58% -20.52% 6.48% 33.87% -27.39%
ALFNet 34.26% 52.63% -18.37% 35.42% 54.73% -19.32% 42.24% 50.00% -7.76%
CSP 40.72% 53.95% -13.23% 36.92% 52.17% -15.25% 38.72% 42.74% -4.02%
MGAN 40.38% $4$35% 33.20% 48.85% -15.65% 30.68% 40.32% -9.65%
PRNet 48.33% 68.42% -20.09% 43.46% 63.17% -19.71% 46.67% 50.81% -4.14%
Average 23.53% 41.61% -18.08% 23.13% 44.18% -21.05% 26.04% 44.35% -18.32%
Skin Tone (LS: Light Skin, DS: Dark Skin)
Detectors level 3 level 2 level 1
LS MR DS MR EOD LS MR DS MR EOD LS MR DS MR EOD
YOLOX 4.15% 4.91% -0.75% 5.75% 3.35% 2.40% 2.52% 2.17% 0.35%
RetinaNet 7.27% 5.28% 1.99% 8.60% 4.18% 4.41% 10.08% 2.17% 7.91%
Faster RCNN 5.04% 4.91% 0.14% 6.16% 2.51% 3.65% 6.72% 2.17% 4.55%
Cascade RCNN 4.01% 4.53% -0.52% 5.64% 2.09% 3.55% 5.04% 4.35% 0.69%
ALFNet 37.09% 36.60% 0.49% 42.62% 45.40% -2.78% 72.27% 71.74% 0.53%
CSP 53.26% 57.74% -4.47% 62.40% 65.48% -3.08% 90.76% 97.83% -7.07%
MGAN 49.26% 50.19% -0.93% 52.30% 54.81% -2.51% 84.03% 82.61% 1.42%
PRNet 54.15% 54.72% -0.56% 59.92% 59.21% 0.71% 86.55% 89.13% -2.58%
Average 26.78% 27.36% -0.58% 30.42% 29.63% 0.80% 44.75% 44.02% 0.73%
Following the roadmap outlined in RQ2.1, we begin by comparing the overall miss rates of pedestrian detectors across different contrast levels, as presented in Table 8. As explained in Section 3.3.3, we categorize driving scenarios into three contrast levels, with a higher level indicating greater contrast. However, we do not observe a consistent pattern in the results. Specifically, concerning light-skin and dark-skin pedestrians, we note a decrease in miss rates overall with increasing contrast. Nonetheless, this pattern is not observed across other demographic groups.
We then examine the shifts in fairness, as evidenced by the trends in EOD with increasing contrast, as outlined in Table 8. Our analysis reveals a consistent pattern: on average, detectors exhibit the most biased results (i.e., the largest absolute value of EOD) under level 2, while demonstrating the fairest outcomes under level 3.
Specifically, for gender, detectors achieve -1.63% EOD under level 2 (most biased), -1.26% under level 1, and 0.93% under level 3 (fairest). For age, detectors achieve -21.05% EOD under level 2 (most biased), -18.32% under level 1, and 18.08% under level 3 (fairest). For skin tone, detectors achieve 0.80% EOD under level 2 (most biased), 0.73% under level 1, and -0.58% under level 3 (fairest).
This finding is further supported by the observation that under level 2, the highest number of detectors exhibit significant biases for gender, age, and skin tone, whereas under level 3, the fewest detectors show significant biases. Specifically, under level 2, five detectors demonstrate significant biases regarding gender, whereas under level 3, only one detector does so. Under level 2, four detectors show significant biases regarding skin tone, whereas under level 3, no detector exhibits significant biases.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false]
Ans. to RQ2.2: We classify driving scenarios into three contrast levels and observe that while there is not a clear pattern in the overall detection performance change with contrast variation, the highest contrast level consistently produces the fairest detection results across gender, age, and skin tone.
@@
As decribed in Section 3.3.3, we consider two weather conditions: non-rainy and rainy.
@@
Table 9 presents the miss rates of eight pedestrian detectors under the two conditions, with statistically significant unfairness results shaded. Overall, the miss rate for each demographic group increases in rainy weather conditions. This escalation could be attributed to droplets covering the camera and disrupting the detectors. A follow-up question is whether rainy weather fairly increase the miss rate for different demographic groups.